### v1

@@ Acknowledgements.
M. E. R. is thankful to the Observatory of Brera-Merate for the kind hospitality during the development of this project, and to Dr. Lara Nava for the insightful discussions that helped shaping this project. This work has been partially supported by the research programme Athena with project number 184.034.002, which is financed by the Dutch Research Council (NWO). This research has made use of data obtained through the High Energy Astrophysics Science Archive Research Center Online Service, provided by the NASA/Goddard Space Flight Center, and specifically this work made use of public Fermi/GBM and Fermi/LAT data.
@@ Appendix A Time-integrated spectrum of GRBs showing a cutoff
Figure 9: Time-integrated spectra of the GRBs showing the presence of an exponential cut-off at high-energies, i.e. best-fitted by the either the SBPLCUTOFF or 2SBPLCUTOFF function (solid line), excluding GRB 140206, whose spectrum has been already reported in Fig. 3. The best fit values of the corresponding models are reported in Table 2. For each plot, data correspond to the signal detected from 10 keV to 10 GeV by the different instruments on board Fermi (colour-coded as shown in the legend). Data have been rebinned for graphical purposes.

### v2

@@ Acknowledgements.
We would like to thank the anonymous referee for carefully reading our manuscript and for providing valuable comments, which improved our manuscript. M. E. R. is thankful to the Observatory of Brera-Merate for the kind hospitality during the development of this project, and to Dr. Lara Nava for the insightful discussions that helped shape this project. This work has been partially supported by the research programme Athena with project number 184.034.002, which is financed by the Dutch Research Council (NWO). This research has made use of data obtained through the High Energy Astrophysics Science Archive Research Center Online Service, provided by the NASA/Goddard Space Flight Center, and specifically, this work made use of public Fermi/GBM and Fermi/LAT data.
@@ Appendix A Time-integrated spectrum of GRBs showing a cutoff
Figure 9: Time-integrated spectra of the GRBs showing the presence of an exponential cutoff at high energies, i.e., those that are best fitted by either the SBPLCUTOFF or 2SBPLCUTOFF function (solid line), excluding GRB 140206, whose spectrum has already been reported in Fig. 3. The best-fit values of the corresponding models are reported in Table 2. For each plot, data correspond to the signal detected from 10 keV to 10 GeV by the different instruments on board Fermi (color-coded as shown in the legend). Data have been rebinned for graphical purposes.